If two moles of Nitrogen is heated from 20 to 30◦C, what are the changes in the translational, rotational and total internal energy?
1
Expert's answer
2020-04-13T09:57:47-0400

As per the given question,

Total number of moles of the Nitrogen gas(n)=2

Initial temperature "(T_1)=20^\circ C=273+30=293K"

Final temperature "(T_2)=30^\circ C=303K"

Boltzmann constant "(k)=1.38\times 10^{-23}J/K"

"R=8.32 mol^{-1}k^{-1}"

Change in transnational "=\dfrac{3R(T_2-T_1)}{2}\times n"

"=3\times 8.32\times (303-293)=249.6J"

Change in rotational energy "=nRT"

"=2\times 8.31\times (303-293)=166.2J"

Change in internal energy="249.6-166.2=83.4J"